# Stochastic Approximation with Block Coordinate Optimal Stepsizes

### v1

@@ 3.3 BCOS with decoupled weight decay
A more effective way in practice is to use decoupled weight decay as proposed in the AdamW algorithm Loshchilov2019AdamW . Specifically, we apply weight decay separately in the BCOS update:
@@
Algorithm 4 shows BCOSW-c, which uses the conditional estimator. Other variants, namely BCOSW-g and BCOSW-m, can be obtained similarly. A PyTorch implementation of all variants of BCOS and BCOSW is given in Appendix A. The code is also available at https://github.com/facebookresearch/bcos.
@@ 4 Numerical experiments
In this section, we present preliminary experiments to compare BCOS with Adam, specifically their variants with decoupled weight decay. Among the BCOSW family, we focus on BCOSW-c (Algorithm 4).
Our first set of experiments is conducted on training a small GPT2 model with 124 million parameters radford2019language on the OpenWebText dataset Gokaslan2019OpenWeb . We use global batch size 512 and run all experiments for 100k iterations. We use linear warmup on the stepsize schedule $\{\alpha_{t}\}$ to reach $\alpha_{\text{max}}$ in the first 2k iterations, and then use cosine decay to toward the final stepsize $\alpha_{\text{min}}=0.01\alpha_{\text{max}}$. The default hyper-parameters are chosen (based on a coarse sweep) as: peak stepsize $\alpha_{\text{max}}=0.002$, $\epsilon=10^{-6}$ and weight decay $\lambda=0.1$.
@@
Figure 3 illustrates the difference between algorithms with and without decoupled weight decay. BCOS-c converges to much higher loss than BCOSW-c, and different values of $\lambda$ (weight decay) make a dramatic difference for BCOS-c but cause little change to BCOSW-c. The same phenomenon happens for Adam versus AdamW, and we again observe more spikes from the loss curves of Adam(W) than BCOS(W)-c.
@@
[…]
@@
These preliminary experiments demonstrate that BCOSW-c can obtain competitive performance compared with the state-of-the-art method AdamW, but with fewer optimizer states and fewer hyper-parameters to tune. More comprehensive empirical studies are needed to fully understand its potential in training large deep learning models.
@@ 5 Convergence analysis
In this section, we present convergence analysis for BCOS with a general search direction and decoupled weight decay (see Section 3.3). To simplify presentation, we stick with the setting of single-coordinate blocks. Specifically, we analyze the algorithm
$x_{t+1}=(1-\alpha_{t}\lambda)x_{t}-\gamma_{t}\odot d_{t},\quad\text{where}\quad\gamma_{t}=\frac{\alpha_{t}}{\sqrt{v_{t}+\epsilon}}.$ (28)
Here, stepsize $\gamma_{t}$ takes a slightly different form from the one used throughout Section 3, where $\epsilon$ is outside the square-root in the denominator. The resulting difference between the two forms is negligible for corresponding choices of $\epsilon$, with the one in (28) being squared (e.g., being $10^{-12}$ if the one in Section 3 is $10^{-6}$). We choose the form in (28) for ease of presenting the analysis.
Recall that $v_{t}\in\mathbf{R}_{+}^{n}$ is an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. Our analysis consists of two stages:
@@
    It is called “conceptual” because we cannot compute $\mathbf{E}_{t}[d_{t}^{2}]$ exactly in practice. But its analysis is essential for understanding and analyzing the practical method. Our analysis relies on an aiming condition, which originates from the classical stochastic approximation literature RobbinsMonro51 ; Blum1954multidim ; chung1954 ; Sakrison1966 . Under this assumption, we establish almost sure convergence of $x_{t}$ to a target point $x_{*}$ and the $\mathcal{O}(1/t)$ convergence rate of $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$.
@@
The rest of this section is organized as follows. In Section 5.1, we introduce the aiming condition and discuss its connection with convexity. We present the convergence analysis for the conceptual BCOS method in Section 5.2 and then the analysis for the practical BCOS method in Section 5.3. Our analysis is very general and applies to the whole BCOS family of methods, including RMSprop, Adam, AdamW, and variants of spectral BCOS. To obtain specific results for each method, we only need to estimate the bias and variance of the second-moment estimator $v_{t}$, which are presented in Section 5.4.
@@ 5.1 Aiming condition
First consider the stochastic approximation method (7). A typical set of assumptions in the classical stochastic approximation literature (specifically Sakrison1966 ) includes:
@@
Under these conditions, the sequence $\{x_{t}\}$ converges to $x_{*}$ in the mean-square sense, i.e., $\lim_{t\to\infty}\mathbf{E}[\|x_{t}-x_{*}\|^{2}]=0$ Sakrison1966 . In particular, the last condition (31) ensures that the expected direction $\mathbf{E}_{\xi}[g(x,\xi)]$ always points towards the target $x_{*}$, implying that the distance from $x_{*}$ can be reduced along this direction (with appropriate stepsize). For this reason, we call it an aiming condition. Similar results can be established under weaker aiming conditions than (31); for example, without the upper bound on the inner product. But they are more subtle and require more sophisticated analysis (e.g., RobbinsMonro51, ; Wolfowitz1952, ; Blum1954prob-one, ; Blum1954multidim, ; Dvoretzky1956, ; DermanSacks1959, ; Wasan69book, ).
Following the same idea, we rewrite the conceptual BCOS method (29) as
@@
Assumption A (Aiming condition).
@@
We make the following remarks regarding Assumption A.
@@
  • We do not need an upper bound on the inner product as in (31) due to the normalization effect of dividing $d_{t}$ by $\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}$ in BCOS. A similar one-sided aiming condition was used in Blum1954multidim to establish almost sure convergence of the first multi-dimensional extension of the Robbins-Monro method.
@@
Intuitively, the inner product in the aiming condition is predominantly determined by the coordinates that are far from optimal, in other words, the coordinates with relatively large $|x_{t,k}-x_{*,k}|$ and having a large SiF $\rho_{t,k}$ (effectively high signal-to-noise ratio).
5.1.1 Connection with convexity
In the context of stochastic optimization, i.e., minimizing $F(x)=\mathbf{E}_{\xi}[f(x,\xi)]$ with $g(x,\xi)=\nabla f(x,\xi)$, the classical aiming condition (31) is implied by $F$ being $\mu$-strongly convex and $L$-smooth (e.g., Nesterov04book, ). But in general it is a much weaker condition. Our aiming condition in Assumption A does not relate directly to smoothness and is not always implied by (strong) convexity either. However, it does have overlapping characteristics with convexity.
This is a direct consequence of convexity, which is characterized by (e.g., Nesterov04book, ):
However, other than the full-dimensional block case, the aiming condition (34) may exhibit significant departure from convexity. As an example, consider the case of single coordinate blocks with uniform SiF, i.e., $\rho_{t,k}=\rho_{t,k^{\prime}}$ for all $k,k^{\prime}\in\{1,\ldots,n\}$. In light of the equivalent formulation using SiF in (33), the corresponding aiming condition becomes
Comparing this condition with convexity, it is easy to construct examples that satisfy one of them while failing the other. We provide some simple examples in Appendix B.1.
@@ 5.2 Analysis of conceptual BCOS
Lemma 5.1.
Suppose Assumption A holds, $\alpha_{t}\geq 0$ and $\alpha_{t}\lambda<1$ for all $t\geq 0$. Then the sequence $\{x_{t}\}$ generated by (29) satisfies, for all $t\geq 0$,
@@
Proof.
First, we notice that the aiming condition (32) is equivalent to
@@
where the last two inequalities are due to Hölder’s inequality $\langle x_{*},\sqrt{\rho_{t}}\rangle\leq\|x_{*}\|_{1}\|\sqrt{\rho_{t}}\|_{\infty}$ and the fact that $\|\sqrt{\rho_{t}}\|_{\infty}\leq 1$, respectively. Plugging the above equality and inequalities into (40) yields the desired result. ∎
As a direct consequence of Lemma 5.1, there exists sufficiently small $\alpha_{t}>0$ such that
@@ 5.2.1 Almost sure convergence
The following theorem establishes almost sure convergence of the conceptual BCOS algorithm under a classical condition on the sequence $\{\alpha_{t}\}$.
Theorem 5.1 (Almost sure convergence).
@@
The under Assumption A, the sequence $\{x_{t}\}$ generated by (29) satisfies $\left\|x_{t}-x_{\ast}\right\|\to 0$ a.s.
The proof of Theorem 5.1 is based on the following almost supermartingale lemma of Robbins and Siegmund ROBBINS1971 .
Lemma 5.2 (“Almost supermartingale”, (ROBBINS1971, , Theorem 1)).
) Let $(\Omega,\mathcal{F},P)$ be a probability space, and $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots$ be a sequence of sub-$\sigma$-algebras of $\mathcal{F}$. For each $t\geq 0$, let $X_{t}$, $a_{t}$, $b_{t}$ and $c_{t}$ be non-negative $\mathcal{F}_{t}$-measurable random variables such that
@@
Proof of Theorem 5.1.
Define $X_{t}:=\left\|x_{t}-x_{\ast}\right\|^{2}$ and $\mathcal{F}_{t}$ to be the $\sigma$-algebra generated by $X_{0},\cdots,X_{t}$. Lemma 5.1 implies the following recursion, in the form of (42),
@@
Therefore, all the assumptions in Lemma 5.2 are satisfied. We may now apply the lemma to conclude that $X_{t}=\left\|x_{t}-x_{\ast}\right\|^{2}\to X$ for some $X<\infty$ and
[…]
This is compatible with the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$ only if $\left\|x_{t}-x_{*}\right\|^{2}\to 0$ a.s. ∎
@@ 5.2.2 Rates of convergence
Corollary 5.1.
Consider the conceptual BCOS method (29) and fix the stepsize schedule $\alpha_{t}=\alpha$ where $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption A, we have
@@
Proof.
Recursive application of Lemma 5.1 with $\alpha_{t}=\alpha$ yields
@@
which is the desired result. ∎
@@
Theorem 5.2.
Consider the algorithm (29) with stepsize schedule $\alpha_{t}=\frac{\alpha}{t+1}$. Suppose the constant $\alpha$ satisfies $11/2<\alpha\lambda<$ 1 and Assumption A holds. Then we have for all $t\geq 0$,
@@
Our proof of Theorem 5.2 is based on Lemma 5.1 and a classical result of Chung chung1954 .
Lemma 5.3 (Chung’s lemma, (chung1954, , Lemma 1)).
@@
Proof of Theorem 5.2.
Taking total expectation on both sides of (37), we obtain
@@
which is in the form of (43) except with the index shift $t\to t+1$. Applying Lemma 5.3 gives the desired result. ∎
Theorem 5.2 requires $\alpha\lambda>1/2$ to obtain the $\mathcal{O}(1/t)$ rate of convergence. We can remove this condition by working with the stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ with $p\in(\frac{1}{2},1)$. In this case we have an asymptotic $\mathcal{O}(1/t^{p})$ rate, as stated in the following theorem.
Theorem 5.3.
Consider the algorithm (29) with stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ where $p\in\left(\frac{1}{2},1\right)$ and the constant $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption A, we have
@@
The proof is based on another lemma of Chung chung1954 , restated below.
Lemma 5.4 ((chung1954, , Lemma 4)).
@@
Proof of Theorem 5.3.
Substituting $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ into (44), we obtain
@@
The desired result follows by applying Lemma 5.4 with the definitions $X_{t+1}=\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]$, $a_{t}=a=2\alpha\lambda$, $b=\alpha^{2}\left(\lambda^{2}\mathbf{E}[\left\|x_{0}-x_{\ast}\right\|^{2}]+(1+K_{p})B_{\ast}\right)$ and $q=2p$. ∎
5.2.3 Analysis for not using decoupled weight decay
Here we briefly explain the convergence analysis of BCOS in the case of $\lambda=0$, or in the case of $\lambda>0$ but without using decoupled weight decay. In other words, we consider

### v2

@@ 3.3 BCOS with decoupled weight decay
A more effective way in practice is to use decoupled weight decay as proposed in the AdamW algorithm (Loshchilov2019AdamW). Specifically, we apply weight decay separately in the BCOS update:
@@
Algorithm 4 shows the BCOSW framework, which can be instantiated with specific $d_{t}$ and $v_{t}$ as in BCOS-g, -m or -c. A PyTorch implementation of all variants is available at
https://github.com/facebookresearch/bcos.
@@ 4 Numerical experiments
In this section, we present preliminary experiments to compare BCOS with Adam, specifically their variants with decoupled weight decay: AdamW and BCOSW. Among the BCOSW family, we focus on BCOSW-c.
Our first set of experiments is conducted on training a small GPT2 model with 124 million parameters (radford2019language) on the OpenWebText dataset (Gokaslan2019OpenWeb). We use global batch size 512 and run all experiments for 100k iterations. We use linear warmup on the stepsize schedule $\{\alpha_{t}\}$ to reach $\alpha_{\text{max}}$ in the first 2k iterations, and then use cosine decay to toward the final stepsize $\alpha_{\text{min}}=0.01\alpha_{\text{max}}$. The default hyper-parameters are chosen (based on a coarse sweep) as: peak stepsize $\alpha_{\text{max}}=0.002$, $\epsilon=10^{-12}$ and weight decay $\lambda=0.1$.
@@
Figure 3 illustrates the difference between algorithms with and without decoupled weight decay. For Adam and BCOS-c, the weight decay is accounted for in the gradient with $g_{t}=\nabla f(x_{t},\xi_{t})+\lambda x_{t}$. BCOS-c converges to much higher loss than BCOSW-c, and different values of $\lambda$ (weight decay) make a dramatic difference for BCOS-c but cause little change to BCOSW-c. The same phenomenon happens for Adam versus AdamW, and we again observe more spikes from the loss curves of Adam(W) than BCOS(W)-c.
@@
Finally, in Figure 5, we compare different algorithms for training ResNet-20 (he2016resnet) on the CIFAR10 dataset (krizhevsky2009learning), and also training the Vision Transformer (ViT) (touvron21vit) on the ImageNet dataset (imagenet2009). For the ResNet task, we tried both cosine decay (drop by factor 100) and step decay (drop by 10 at epochs 80, 120, 150). The hyper-parameters chosen are: $\beta=0.9$ for SGD and BCOSW-c, and $\beta_{1,2}=(0.9,\,0.99)$ for AdamW. We observe that the best-performing stepsize schedules are quite different for different methods. This prompts the need to tune hyper-parameters for BCOSW for different tasks even though it shares similar tuned hyper-parameters as AdamW on the GPT2 task.
@@
These preliminary experiments demonstrate that BCOSW-c can obtain competitive performance compared with the state-of-the-art method AdamW, but require less memory and fewer hyper-parameters to tune. More comprehensive empirical studies are needed to fully understand its potential in training larger deep learning models.
@@ 5 Convergence analysis
In this section, we present convergence analysis for BCOS(W). To simplify presentation, we stick with the setting of single-coordinate blocks. Specifically, we analyze the algorithm
$x_{t+1}=(1-\alpha_{t}\lambda)x_{t}-\gamma_{t}\odot d_{t},\quad\text{where}\quad\gamma_{t}=\frac{\alpha_{t}}{\sqrt{v_{t}+\epsilon}},$ (29)
where $d_{t}$ is a general search direction and $v_{t}\in\mathbf{R}_{+}^{n}$ is an online estimator of $\mathbf{E}_{t}[d_{t}^{2}]$. Our analysis consists of two stages:
@@
    It is called “conceptual” because we cannot compute $\mathbf{E}_{t}[d_{t}^{2}]$ exactly in practice. But its analysis is essential for understanding and analyzing the practical method. Our analysis relies on an aiming condition, which originates from the classical stochastic approximation literature (RobbinsMonro51; Blum1954multidim; chung1954; Sakrison1966). Under this assumption, we establish almost sure convergence of $x_{t}$ to a target point $x_{*}$ and the $\mathcal{O}(1/t)$ convergence rate of $\mathbf{E}[\|x_{t}-x_{*}\|^{2}]$.
@@
The rest of this section is organized as follows. In Section 5.1, we introduce the aiming condition and give some empirical evidence. We present the convergence analysis for the conceptual BCOS method in Section 5.2 and then the analysis for the practical BCOS method in Section 5.3. Our analysis is very general and applies to the whole BCOS family of methods, including the sign-gradient method, RMSprop and Adam(W). To obtain specific results for each method, we only need to estimate the bias and variance of the second-moment estimator $v_{t}$, which are presented in Section 5.4.
@@ 5.1 Aiming condition
First consider the stochastic approximation method (7). A typical set of assumptions in the classical stochastic approximation literature includes:
@@
Under these conditions, the sequence $\{x_{t}\}$ converges to $x_{*}$ in the mean-square sense, i.e., $\lim_{t\to\infty}\mathbf{E}[\|x_{t}-x_{*}\|^{2}]=0$ (e.g., Sakrison1966). In particular, the last condition (32) ensures that the direction $\mathbf{E}_{\xi}[g(x,\xi)]$ always points towards the target $x_{*}$, implying that the distance from $x_{*}$ can be reduced along this direction (with appropriate stepsize). For this reason, we call it an aiming condition.
In the context of stochastic optimization, i.e., minimizing $F(x)=\mathbf{E}_{\xi}[f(x,\xi)]$ with $g(x,\xi)=\nabla f(x,\xi)$ and $\mathbf{E}_{\xi}[g(x,\xi)]=\nabla F(x)$, the aiming condition (32) is implied by $F$ being $\mu$-strongly convex and $L$-smooth (e.g., Nesterov04book). But in general it is a much weaker condition. Moreover, the upper bound on the inner product can often be removed by leveraging more sophisticated analysis (e.g., RobbinsMonro51; Wolfowitz1952) or careful choice of the search direction $g(x,\xi)$ (e.g. Blum1954multidim; Dvoretzky1956). This is the case of our aiming condition, which we present next.
In order to allow reduction of the expected distance from $x_{t+1}$ to $x_{*}$, we rewrite the conceptual BCOS method (30) as
@@
Assumption 1 (Aiming condition)
@@
We make the following remarks regarding Assumption 1.
@@
  • We do not need an upper bound on the inner product as in (32) due to the normalization effect of dividing $d_{t}$ by $\sqrt{\mathbf{E}_{t}[d_{t}^{2}]}$ in BCOS. A similar one-sided aiming condition was used in (Blum1954multidim) to establish almost sure convergence of the first multi-dimensional extension of the Robbins-Monro method.
@@
Intuitively, the inner product in the aiming condition is predominantly determined by the coordinates that are far from target, in other words, the coordinates with relatively large $|x_{t,k}-x_{*,k}|$ and having a large SiF $\rho_{t,k}$ (effectively high signal-to-noise ratio).
In general we cannot check the aiming condition analytically, since it involves conditional expectations and an unknown target point $x_{*}$. However, we can check empirically if the inner product $\langle x_{t}-x_{*},d_{t}/\sqrt{v_{t}}+\lambda x_{t}\rangle$ is positive while running any particular algorithm. In order to do so, we approximate $x_{*}$ with the last iterate after running the algorithm for many iterations. Figure 6 and 7 depict the empirical aiming condition while running SGD, SGD with momentum, BCOS-c and BCOSW-c on GPT2-124M with hyper-parameters as chosen in Section 4. Specifically, we record the cosines between $x_{t}-x_{\ast}$ and various directions including stochastic gradients $g_{t}$, momentum $m_{t}$, $m_{t}/\sqrt{v_{t}}$ and $m_{t}/\sqrt{v_{t}}+\lambda x_{t}$ along the training trajectory. We notice that the cosines of the angles between $x_{t}-x_{*}$ and the search directions are positive (correct aiming) most of the time, especially in the later stage of training. Moreover, the actual update directions in each algorithm exhibit the highest cosine values with $x_{t}-x_{\ast}$ among all these directions.
The above results give empirical evidence that the aiming condition can be much weaker than convexity and often holds in training deep learning models. We explain the connection between aiming and convexity in more detail in Appendix A.1.
Figure 6: Empirical aiming conditions on GPT2. Left: SGD. Right: SGD with momentum.
Figure 7: Empirical aiming conditions on GPT2. Left: BCOS-c. Right: BCOSW-c.
@@ 5.2 Analysis of conceptual BCOS
Lemma 1
Suppose Assumption 1 holds, $\alpha_{t}\geq 0$ and $\alpha_{t}\lambda<1$ for all $t\geq 0$. Then the sequence $\{x_{t}\}$ generated by (30) satisfies, for all $t\geq 0$,
@@
Proof First, we notice that the aiming condition (33) is equivalent to
@@
where the last two inequalities are due to Hölder’s inequality $\langle x_{*},\sqrt{\rho_{t}}\rangle\leq\|x_{*}\|_{1}\|\sqrt{\rho_{t}}\|_{\infty}$ and the fact that $\|\sqrt{\rho_{t}}\|_{\infty}\leq 1$, respectively. Plugging the above equality and inequalities into (38) yields the desired result.
As a direct consequence of Lemma 1, there exists sufficiently small $\alpha_{t}>0$ such that
@@ 5.2.1 Almost sure convergence
The following theorem establishes almost sure (a.s.) convergence of the conceptual BCOS algorithm under a classical condition on the sequence $\{\alpha_{t}\}$.
Theorem 2 (Almost sure convergence)
@@
The under Assumption 1, the sequence $\{x_{t}\}$ generated by (30) satisfies $\left\|x_{t}-x_{\ast}\right\|\to 0$ a.s.
The proof of Theorem 2 is based on the following almost supermartingale lemma of ROBBINS1971.
Lemma 3 (ROBBINS1971)
Let $(\Omega,\mathcal{F},P)$ be a probability space, and $\mathcal{F}_{0}\subset\mathcal{F}_{1}\subset\ldots$ be a sequence of sub-$\sigma$-algebras of $\mathcal{F}$. For each $t\geq 0$, let $X_{t}$, $a_{t}$, $b_{t}$ and $c_{t}$ be non-negative $\mathcal{F}_{t}$-measurable random variables such that
@@
Proof [Proof of Theorem 2] Define $X_{t}:=\left\|x_{t}-x_{\ast}\right\|^{2}$ and $\mathcal{F}_{t}$ to be the $\sigma$-algebra generated by $X_{0},\cdots,X_{t}$. Lemma 1 implies the following recursion, in the form of (40),
@@
Therefore, all the assumptions in Lemma 3 are satisfied. We may now apply the lemma to conclude that $X_{t}=\left\|x_{t}-x_{\ast}\right\|^{2}\to X$ for some $X<\infty$ and
$\sum_{t=0}^{\infty}c_{t}=\sum_{t=0}^{\infty}2\alpha_{t}\lambda\left\|x_{t}-x_{\ast}\right\|^{2}<\infty\quad\text{ a.s.}$
This is compatible with the assumption $\sum_{t=0}^{\infty}\alpha_{t}=\infty$ only if $\left\|x_{t}-x_{*}\right\|^{2}\to 0$ a.s.
@@ 5.2.2 Rates of convergence
Corollary 4
Consider the conceptual BCOS method (30) with a constant step size schedule $\alpha_{t}=\alpha$ where $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption 1, we have
@@
Proof Recursive application of Lemma 1 with $\alpha_{t}=\alpha$ and taking total expectation yields
@@
which is the desired result.
@@
Theorem 5
Consider the algorithm (30) with stepsize schedule $\alpha_{t}=\frac{\alpha}{t+1}$. Suppose the constant $\alpha$ satisfies $1/2<\alpha\lambda<1$ and Assumption 1 holds. Then we have for all $t\geq 0$,
@@
Our proof of Theorem 5 is based on Lemma 1 and a classical result of chung1954.
Lemma 6 (chung1954)
@@
Proof [Proof of Theorem 5] Taking total expectation on both sides of (35), we obtain
@@
which is in the form of (41) except with the index shift $t\to t+1$. Applying Lemma 6 gives the desired result.
Theorem 5 requires $\alpha\lambda>1/2$ to obtain the $\mathcal{O}(1/t)$ rate of convergence. We can remove this condition by working with the stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ with $p\in(\frac{1}{2},1)$. In this case we have an asymptotic $\mathcal{O}(1/t^{p})$ rate, as stated in the following theorem.
Theorem 7
Consider the algorithm (30) with stepsize schedule $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ where $p\in\left(\frac{1}{2},1\right)$ and the constant $\alpha$ satisfies $\alpha\lambda<1$. Then under Assumption 1, we have
@@
The proof is based on another lemma of chung1954, restated below.
Lemma 8 (chung1954)
@@
Proof [Proof of Theorem 7] Substituting $\alpha_{t}=\frac{\alpha}{(t+1)^{p}}$ into (42), we obtain
@@
The desired result follows by applying Lemma 8 with the definitions $X_{t+1}=\mathbf{E}[\left\|x_{t}-x_{\ast}\right\|^{2}]$, $a_{t}=a=2\alpha\lambda$, $b=\alpha^{2}\left(\lambda^{2}\mathbf{E}[\left\|x_{0}-x_{\ast}\right\|^{2}]+(1+K_{p})B_{\ast}\right)$ and $q=2p$.
5.2.3 Convergence analysis with $\lambda=0$
Here we briefly explain the convergence analysis of BCOS in the case of $\lambda=0$. In other words, we consider